OpenStudy (anonymous):

how do i solve this question? please help me 1) y=4(x-1)*2+2, 2) y=5(x-2)*2+15, 3) y=-2(x+2)*2-1, 4) y=3(x-6)*2-11, 5)y= (x-10)*2-25

OpenStudy (azureilai):

You use the distributive property and the combine like terms. For example, on the first one y=4(x-1)*2+2 =(4x-4)*2+2 =8x-8+2 =8x-6 Do the same for the others

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

please i need more answer? please help me

OpenStudy (azureilai):

Just look at my example and apply it to the other equations. Or you can do them and I can check the answers for you.